OpenStudy (anonymous):

Simplify 6y^-1 . 2x^6 . 4v^-5 x^8 yv^7 use on positive exponents in your answer

OpenStudy (jdoe0001):

keeping in mind that \(\bf a^{-n} = \cfrac{1}{a^n}\) then \(\bf 6y^{-1}\cdot 2x^6 \cdot 4v^{-5} x^8 yv^7\\ \quad \\ \quad \\ 6y^{-1}\implies 6 \cdot \cfrac{1}{y^1}\implies \cfrac{6}{y}\\ \quad \\ 4v^{-5}\implies 4 \cdot \cfrac{1}{v^5}\implies \cfrac{4}{v^5}\qquad thus\\ \quad \\ 6y^{-1}\cdot 2x^6 \cdot 4v^{-5} x^8 yv^7\implies \cfrac{6}{y}\cdot 2x^6\cdot \cfrac{4}{v^5}\cdot x^8 yv^7\)